A train leaves a station and travels east at 72 km/h. Three hours later a second train leaves on a parrallel track and travels east at 120 km/h. When will it overtake the first train?
Answer by josmiceli(6796)

It usually helps to make a chart for these problems
*
*
-----------------------------------------------------------------------
...............distance travelled.....|.rate km/hr.|..time hrs
-----------------------------------------------------------------------
slow train........72t....................|.........72..........|......t
-----------------------------------------------------------------------
fast train........120(t - 3).........|........120..........|......t - 3
-----------------------------------------------------------------------
The time the fast train is running is 3 hours less than the slow one,
but the distance it travels is the same as the fast one. That's where
they meet. So set the distances in the chart equal to eachother.
72t = 120(t - 3)
72t = 120t - 360
subtract 72t from each side
48t = 360
t = 7.5
This is the time it takes the slow train to get to where they meet,
but the problem wants ti know weh the fast train will overtake it
It will take the fast train (t - 3) hours to overtake the slow one.
t = 7.5
(t - 3) = (7.5 - 3) = 4.5 hours
check the answer
72t = 120(t - 3)
72(7.5) = 120(7.5 - 3)
540 = 120*4.5
540 = 540
OK